OpenStudy (anonymous):

a right triangle has a hypotenuse of 23 and a length of 31 what is the length of side x round your answer if necessary

OpenStudy (rane):

use Pythagorus theorem c^2=a^2+b^2

OpenStudy (anonymous):

how do i know which one goes to which?

OpenStudy (rane):

c is always the hypotenuse

OpenStudy (rane):

a is the side length

OpenStudy (jdoe0001):

the LONGEST one is always hypotenuse

OpenStudy (anonymous):

okay so how do i set it up

OpenStudy (jdoe0001):

now which is "a" or "b", they're interchangeable for the use of the pythagorean theorem

OpenStudy (anonymous):

i dont understand

OpenStudy (jdoe0001):

as @RANE said -> c^2=a^2+b^2 <-, use the LONGEST value for "c" c=hypotenuse

OpenStudy (rane):

ust subsitute the values in the formula c=23 a=31 b=? so, 23^2= 31^2+b^2 529=961+b^2 b^2= 961-529 = \[\sqrt{432}\] = 20.8

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

If ABC XYZ, what is the length of ?

OpenStudy (rane):

???